OpenStudy (anonymous):

The problem: \[y=3x^2 + sin (2x)\]; write down a 2nd order differential equation involving only d^2x/dy^2, functions of x and y, and constants. Is this just asking for the second differential of the function i.e. \[6-4sin(2x)\]?

OpenStudy (kinggeorge):

My understanding is that it wants you to find an equation relating \(y=3x^2+\sin(2x)\) and \(y''=6-4\sin(2x)\). This equation should ideally be in the form\[f(x)y''+g(x)y=h(x)\] where \(f,g,h\) are functions of \(x\).

OpenStudy (kinggeorge):

It would be even better if you could manipulate the functions so that \(h(x)=0\).